OpenStudy (anonymous):

An astronaut weighs 800 N on the surface of earth. What is the weight of the astronaut 6.37 × 106 m above the surface of the earth? Assume the radius of earth is 6,378 km. 0.0 N 200 N 1,600 N 3,200 N What is the answer and how is it found? Please explain the formula too?:)

OpenStudy (anonymous):

well, you know that \[F=G \frac{ m1xm2 }{ d ^{2} }\] m1=mass of object m2= mass of earth G= Gravitational constant = 6.67300 × 10-11 [m3] [kg-1] [s-2] d= distance from the center of the earth to the object sol. 1. You know that the astronaut is 6.37 × 10^6 m above the surface of the earth. You sum that number to the radius of the earth 6,378 km. (Note: you have to convert that to m, that is 6378x10^6m.) This sum will give you "d". (d= 1275x10^7m.) 2. The mass of Earth m2 = 5.97219 × 10^24 kilograms. The mass of the astronaut is his weight difided by the force of gravity on the surface of the earth (9,8 m/s^2). This is m1= 81.63kg. (This is how much mass he has on earth) 3. You have "d", "m1", "m2" and G is a constante. You can now sub in the formula and get F. This is the force the earth attracts object (Sorry for the bad english, not my first language XD) Anywho. . This is also his weight. That is 200N. PS: Sorry for sending it 3 times. . i had a error.